OpenStudy (anonymous):

Two similar solids are shown below. What is the volume of the larger solid? 5400 ft3 7400 ft3 8200 ft3 6200 ft3

jimthompson5910 (jim_thompson5910):

Drawing is missing

OpenStudy (anonymous):

I keep getting the wrong answer.

jimthompson5910 (jim_thompson5910):

We can say (3^3)/(6^3) = 675/x which becomes 27/216 = 675/x Tell me what you get when you solve for x

OpenStudy (anonymous):

5400

OpenStudy (anonymous):

Thanks!!!

jimthompson5910 (jim_thompson5910):

you got it, 5400 is correct